You need $150,000 in 18 years for your investment needs. If you can earn 6% per year, how much do you need to invest today?

Select one:

a. $62,551.57

b. $52,551.57

c. $72,851.57

d. $55,551.57
Answer #1

b. $52,551.57

Compound interest formula

r. nt A = P(1 +-) n

A = Final amt = 150,000

P = Principle / Present value = ?

r = Interesr rate = 6 % = 0.06

n = no. of times interest applied per time period = 1 (assuming compunded annualy)

t = no. of time periods = 18

0.06 150,000 = P(1+18 1

150,000 = P(1 + 0.0618

P= = 150,000 (1 + 0.06.18

150,000 P= 2.8543

P= 52,551.57